OpenStudy (anonymous):

Evaluate arccos(sin(9pi/4))

OpenStudy (anonymous):

sin(9pi/4)= sin(9pi/4-2*pi) (becoz sin function is periodic) = sin(pi/4)= sqrt(2)/2 arccos(sqrt(2)/2)=pi/4 (becoz cos(pi/4)=sin(pi/4)=sqrt(2)/2)